OpenStudy (anonymous):

If f(x) = 5x - 3, then what is f-1(x) ? Choose one answer. a. x plus 5, over 3 b. x minus 5, over 3 c. x plus 3, over 5 d. x minus 3, over 5

OpenStudy (cwtan):

u mean inverse function? let y=f(x) 5x-3=y 5x=y+3 x=(y+3)/5 so is plus 3 over 5